A toy rocket, mass 1.0 kg, blasts up at a 45o angle from ground level with a kinetic energy of 37 J. To what maximum vertical height does it rise? Round to the nearest tenth of a meter. (Hint: what portion of KE will convert into Ug?)
1
Expert's answer
2020-01-14T09:18:36-0500

The kinetic energy of a rocket

"{\rm KE}=\frac{mv^2}{2}"

So, the initial speed of a rocket


"v=\sqrt{2{\rm KE}/m}=\sqrt{2\times 37/1.0}=8.6\:\rm m/s"

The vertical component of the initial speed


"v_y=v\sin\theta=8.6\times \sin 45^{\circ}=6.1\:\rm m/s"

The maximum vertical height


"h=\frac{v_y^2}{2g}=\frac{6.1^2}{2\times 9.8}=1.9\:\rm m"

Answer: "1.9\:\rm m"